OpenStudy (anonymous):

Find MR when p=√(800-4Q)

OpenStudy (goformit100):

what is the full-form of MR ?

OpenStudy (anonymous):

mentally retard just kidding

OpenStudy (nurali):

Marginal revenue

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

it is just the derivative of that function

OpenStudy (anonymous):

TR=PQ TR=(800-4Q)^(1/2)Q MR=TR´=1/2(800-4Q)^(-1/2)+(800-4Q)^(1/2) =-Q(800-4Q)^-(1/2)+(800-4Q)^(-1/2) = I dont know